OpenStudy (anonymous):

what would the answer for the simplified of (5x)^4

OpenStudy (across):

\((5x)^4=5^4x^4=?\)

OpenStudy (anonymous):

um did u want to know which one it is if so its the first one

OpenStudy (anonymous):

Here's your answer.\[625x ^{4}\] Hope I helped.

OpenStudy (anonymous):

how is it 625x^4